# QLLM: Accurate and Efficient Low-Bitwidth Quantization for Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated unparalleled efficacy in natural language processing. However, their high computational demands and memory overheads hinder their broad deployment. To address this, two quantization strategies emerge, including Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). For LLMs, the billions of parameters make the QAT impractical due to the prohibitive training cost and thus PTQ becomes more prevalent. In existing studies, activation outliers in particular channels are identified as the biggest challenge to PTQ accuracy. They propose to transform the magnitudes from activations to weights, which however offers limited alleviation or suffers from unstable gradients, resulting in a severe performance drop at low-bitwidth. In this paper, we propose QLLM, an accurate and efficient low-bitwidth PTQ method designed for LLMs. QLLM introduces an adaptive channel reassembly technique that reallocates the magnitude of outliers to other channels, thereby mitigating their impact on the quantization range. This is achieved by channel disassembly and channel assembly, which first breaks down the outlier channels into several sub-channels to ensure a more balanced distribution of activation magnitudes. Then similar channels are merged to maintain the original channel number for efficiency. Additionally, an adaptive strategy is designed to autonomously determine the optimal number of sub-channels for channel disassembly. To further compensate for the performance loss caused by quantization, we propose an efficient tuning method that only learns a small number of low-rank weights while freezing the pre-trained quantized model. After training, these low-rank parameters can be fused into the frozen weights without affecting inference. Extensive experiments on LLaMA-1 and LLaMA-2 show that QLLM is able to obtain accurate quantized models efficiently. For example, QLLM quantizes the 4-bit LLaMA-2-70B within 10 hours on a single A100-80G GPU, outperforming the previous state-of-the-art method by 7.89% on the average accuracy across five zero-shot tasks.
@@ 5 Experiments
Quantization settings. In alignment with prior research (Dettmers et al., 2022; Shao et al., 2023), we use per-channel weight quantization and per-token activation quantization. Following (Shao et al., 2023; Liu et al., 2023), we quantize all weights and intermediate activations, with the exception of the Softmax output probability, which is maintained at full precision. Following OmniQuant (Shao et al., 2023), we focus on 4- and 6-bit weights and activations quantization. We exclude 8-bit quantization as SmoothQuant (Xiao et al., 2023) is able to achieve lossless performance.
@@
Implementation details. Following OmniQuant (Shao et al., 2023), we construct the calibration set with 128 randomly sampled sequences from WikiText2, each with a sequence length of 2048. QLLM begins by applying channel reassembly prior to all linear projection layers, excluding the attention output projection layer, followed by performing error correction on the resulting model. The rank of the introduced low-rank parameters is set to 4, and these parameters are trained for 10 epochs with a mini-batch size of 1. We carry out the reconstruction using 4 Attention-FFN blocks. AdamW (Loshchilov & Hutter, 2019) with a linear learning rate decay scheduler is used following (Yao et al., 2022). The learning rate is set to . All training experiments are conducted on a single NVIDIA A100 80G GPU. We use the Language Model Evaluation Harness toolbox (Gao et al., 2021) for evaluation.
@@ 5.1 Main Results
We report the results on LLaMA-1 and LLaMA-2 families in Table 1, and Table A in Appendix. Note that W6A6 has limited hardware support in real-world applications. However, our QLLM still demonstrates performance benefits in these settings, consistently surpassing OmniQuant in terms of lower perplexity across all models on both WikiText2 and C4 and achieving comparable accuracy on 5 zero-shot tasks. While the absolute performance gains with 6-bit quantization might seem modest, this is partly due to the less pronounced effect of activation outliers at this bitwidth. When focusing on extremely low-bitwidth quantization (i.e., 4-bit), activation outliers serve as the performance bottleneck, thereby highlighting the importance of suppressing the outliers. In this case, our QLLM achieves significantly higher zero-shot accuracy and much lower perplexity than the contenders. For example, QLLM quantized 4-bit LLaMA-1-65B outperforms OmniQuant counterpart by an average of 3.42% in accuracy across five zero-shot tasks. Remarkably, for LLaMA-7B, our QLLM even surpasses the QAT method, LLM-QAT + SQ, by 8.6% on the average accuracy, which strongly demonstrates the efficacy of our QLLM.
5.2 Ablation Studies
@@
| LLaMA-1-7B | W6A6 | SQ | 76.65 | 53.11 | 40.10 | 71.52 | 61.88 | 60.65 |
| W6A6 | OS+ | 76.82 | 51.35 | 41.13 | 71.42 | 65.98 | 61.34 | |
| W6A6 | CR | 76.88 | 52.31 | 40.87 | 71.37 | 64.33 | 61.15 | |
[…]
| W4A4 | OS+ | 62.73 | 39.98 | 30.29 | 44.39 | 52.96 | 46.07 | |
| W4A4 | CR | 66.92 | 42.55 | 32.34 | 54.31 | 50.04 | 49.23 | |
| LLaMA-1-13B | W6A6 | SQ | 77.80 | 56.36 | 42.58 | 75.11 | 68.11 | 63.99 |
| W6A6 | OS+ | 78.29 | 56.90 | 43.09 | 75.09 | 69.22 | 64.52 | |
| W6A6 | CR | 78.02 | 56.69 | 42.41 | 74.70 | 70.01 | 64.37 | |
[…]
| W4A4 | OS+ | 63.00 | 40.32 | 30.38 | 53.61 | 51.54 | 47.77 | |
| W4A4 | CR | 67.57 | 43.77 | 31.48 | 60.78 | 56.04 | 51.93 |
Channel reassembly vs. other outlier handling methods. To further show the effectiveness of channel reassembly, we compare our method with previous outlier handling methods which employ gradient-free methods to learn mathematically equivalent transformations. For fair comparisons, we do not apply efficient error correction. From Table 4, all methods exhibit comparable performance at 6-bit quantization. However, for 4-bit quantization, channel reassembly significantly surpasses other methods by a large margin, particularly for larger models.
@@
Effect of efficient gradient-based error correction. After channel reassembly, we implement our QLLM to produce 4-bit LLaMA-7B models with our efficient gradient-based error correction (EEC) and tuning quantized weights directly (TQW) outlined in Section 4.2 to further improve the performance of quantized LLMs and show the results in Table 5. Compared with TQW which tunes all quantized weights, EEC focuses on learning a small set of low-rank weights, which significantly reduces training costs and GPU memory usage while delivering comparable performance. Moreover, the reduced GPU memory demand allows EEC to quantize LLaMA-1-65B on a single 24GB consumer-grade GPU, such as the NVIDIA RTX 4090, a task that is not feasible with TQW. Due to the page limited, we put more results in Section K of the supplementary material.
Inference efficiency. To assess the inference efficiency of our channel reassembly technique, we measure the inference speed of QLLM on NVIDIA RTX 3090 GPUs. We employ W4A4 kernels from QUIK (Ashkboos et al., 2023) codebase. We also conduct a comparative analysis using weight quantization only, utilizing CUDA kernels from AutoGPTQ111https://github.com/PanQiWei/AutoGPTQ. As shown in Table 3, our 4-bit QLLM only incurs 4% additional cost relative to W4A4 but achieves a notable 1.96 speedup over FP16. Notably, our channel reassembly strategy substantially mitigates losses attributed to quantizing outliers (see Table 4), with only a slight extra computational overhead. For the detailed inference cost of channel disassembly and assembly, please refer to Section M of the supplementary material.
@@ 6 Conclusion and Future Work
In this paper, we have proposed an accurate and efficient post-training quantization approach for low-bit LLMs, dubbed QLLM. The core of our QLLM lies in a novel adaptive channel reassembly paradigm that effectively addresses activation outliers, a pivotal factor contributing to the performance bottleneck in quantizing LLMs. The key idea involves reallocating outlier magnitudes to other channels, accomplished through a process of channel disassembly followed by assembly. We have further proposed a quantization-aware, parameter-efficient fine-tuning strategy that leverages calibration data to compensate for the information loss resulting from quantization. Extensive experiments on LLaMA model series have demonstrated the promising performance and training efficiency of our QLLM. In terms of limitations, our proposed channel reassembly involves introducing additional operations to decompose and aggregate channels during runtime, thereby incurring additional inference costs. A potential solution to improve inference efficiency is to explore kernel fusing (Wang et al., 2010), aiming to fuse disassembly, assembly and layer normalization into a single operator. Another way is to aggregate more similar or unimportant channels (Sun et al., 2023) than those disassembled to achieve higher speedup.
@@ Acknowledgments
We sincerely thank Shengwu Jiang for his help in implementing the efficient Triton kernel.
@@ Appendix E More results on LLaMA-2 family
We provide additional results for the LLaMA-2 family in Table A. The observations from these results are consistent with the phenomena identified in the LLaMA-1 family. Note that the varied performance of OmniQuant on W6A6 and W4A4 for LLaMA-2-70B can be attributed to the architecture of LLaMA-2-70B, which employs grouped-query attention (Ainslie et al., 2023) where each group of queries shares a single key and value head. Such architecture makes the learnable equivalent transformation in OmniQuant incompatible with grouped-query attention. In W6A6 settings, the impact of activation outliers is relatively minor, enabling partial learnable equivalent transformations to suffice in maintaining performance. However, in the W4A4 settings, the effect of activation outliers becomes more prominent. Under these conditions, the partial learnable equivalent transformation is insufficient to address the outlier issue, leading to notably poorer performance. Notably, our QLLM significantly outperforms the state-of-the-art post-training quantization (PTQ) methods, demonstrating a substantial margin of improvement in 4-bit quantization. For example, QLLM quantized 4-bit LLaMA-2-65B outperforms SmoothQuant counterpart by an average of 7.89% on the accuracy, which shows the promising results of our method.
@@ Appendix H More results in terms of channel disassembly only
Appendix I More results in terms of the efficient error correction only
Using EEC only without our channel reassembly results in suboptimal performance as it suffers from activation outlier issues. To demonstrate this, we applied EEC only to quantize LLaMA-1-7B to 4-bit, using the same training settings as our QLLM but with varying numbers of calibration samples. From Table E, even with an increased amount of calibration data, the performance of the EEC only significantly lags behind our QLLM. These results strongly demonstrate the effectiveness of channel reassembly in addressing activation outliers, thereby substantially improving performance.
@@ Appendix I More results in terms of the efficient error correction only
Appendix J More results in terms of tuning quantized weights only
The effectiveness of TQW is highly dependent on our channel reassembly. To demonstrate this, we applied TQW only to quantize LLaMA-1-7B to 4-bit using the same training settings as QLLM and show the results in Table F. The results clearly indicate that the absence of our adaptive channel reassembly results in significantly reduced performance for TQW. This underscores the vital role of channel reassembly in addressing activation outliers and thus improving model performance.
@@
Appendix K More comparisons between efficient error correction and tuning quantized weights directly
To further show the effectiveness of our efficient error correction (EEC), we conduct more comparisons between EEC and tuning quantized weights (TQW) directly on small model and report the results in Table G. The results show that employing EEC not only maintains comparable performance but also markedly improves training speed and significantly reduces GPU memory usage over TQW. It is worth noting that there is a trade-off between GPU memory (i.e., #Attn-FFN blocks) and performance. Leveraging EEC even allows us to perform reconstruction for 16 Attention-FFN blocks simultaneously, thereby significantly improving performance while preserving a similar training speed and a reasonable increase in GPU memory.
@@
Appendix L Effect of the weight merging in efficient error correction
As explained in Section 4.2, for , the low-rank weights and bring not only additional inference overhead due to the matrix multiplication between the full-precision and but also extra storage burden. To address this, we perform weight merging by after the reconstruction, which effectively avoids overhead but introduces additional quantization error. For 4-bit quantization, we empirically observe that merging the low-rank weights into the frozen weights using does not lead to an increase in outliers. This finding is supported by the notably low MSE levels for channel-wise , , and maximum/minimum values before and after the weight merging process across in Table H. Moreover, our weight merging only leads to small quantization error, as shown in Table I. Note that even small deviations can aggregate throughout the network, leading to the performance drop. To address this, as shown in Section 4.2, we further employ sequential reconstruction to mitigate errors from previous layers, resulting in only a negligible performance drop. To demonstrate this, we compare the performance of QLLM with and without the weight merging. From Table J, the weight merging only leads to a slight increase in perplexity.
@@
Appendix M More results regarding inference efficiency
Following (Guo et al., 2020; Wang et al., 2020), we use Bit-Operation (BOP) count to measure the theoretical inference complexity of our QLLM. From Table K, our 8-bit QLLM incurs only a marginal increase in BOPs when compared to the INT8 model but substantially lower than those of the FP16 counterpart, which shows the efficiency of our method.
@@
We further show the inference time of channel disassembly and assembly of our QLLM in Table L. From the results, channel disassembly results in additional inference costs due to the extra channels. These additional channels often don’t align with GPU-friendly multiples like 32 or 64, leading to less efficient GPU use. Using our channel assembly maintains the original channel count, ensuring better GPU utilization and mitigating the extra inference costs from disassembly. As a result, the quantized models with both channel disassembly and assembly achieve higher throughput compared to the ones with disassembly only, which demonstrates the necessity of channel assembly.
@@
Appendix N More results regarding training efficiency
We assess the training efficiency of our method in comparison to OmniQuant on a single NVIDIA A100 80G GPU. The GPU training hours for both methods are presented in Table M. The results reveal that the training cost of our QLLM can be up to 1.93 faster than OmniQuant, showing the exceptional training efficiency of our QLLM.
@@
Appendix O Effect of different calibration sets
We apply QLLM to yield 4-bit LLaMA-7B using different calibration sets and report the results in Table N. From the results, we observe that the choices of calibration set have a minor effect, as the performance remains relatively consistent across different sets. For fair comparisons, following OmniQuant (Shao et al., 2023), we use WikiText2 as a calibration set by default.
@@
Appendix P Effect of different numbers of calibration samples
We apply QLLM to yield 4-bit LLaMA-7B using different numbers of calibration samples from WikiText2 and show the results in Table O. The results reveal a positive correlation between the performance of the quantized model and the number of calibration samples, indicating that utilizing more samples generally leads to better performance. This trend underscores the importance of the calibration phase, where leveraging a larger sample pool can provide a more comprehensive representation of the data distribution, enabling more accurate quantization. However, it is also imperative to consider the computational and memory overhead associated with an increasing number of calibration samples. There is an inherent trade-off between achieving higher model performance and maintaining computational efficiency.
@@
Appendix Q More results about the expansion ratios of the quantized LLM

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated unparalleled efficacy in natural language processing. However, their high computational demands and memory overheads hinder their broad deployment. To address this, two quantization strategies emerge, including Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). For LLMs, the billions of parameters make the QAT impractical due to the prohibitive training cost and thus PTQ becomes more prevalent. In existing studies, activation outliers in particular channels are identified as the biggest challenge to PTQ accuracy. They propose to transform the magnitudes from activations to weights, which however offers limited alleviation or suffers from unstable gradients, resulting in a severe performance drop at low-bitwidth. In this paper, we propose QLLM, an accurate and efficient low-bitwidth PTQ method designed for LLMs. QLLM introduces an adaptive channel reassembly technique that reallocates the magnitude of outliers to other channels, thereby mitigating their impact on the quantization range. This is achieved by channel disassembly and channel assembly, which first breaks down the outlier channels into several sub-channels to ensure a more balanced distribution of activation magnitudes. Then similar channels are merged to maintain the original channel number for efficiency. Additionally, an adaptive strategy is designed to autonomously determine the optimal number of sub-channels for channel disassembly. To further compensate for the performance loss caused by quantization, we propose an efficient tuning method that only learns a small number of low-rank weights while freezing the pre-trained quantized model. After training, these low-rank parameters can be fused into the frozen weights without affecting inference. Extensive experiments on LLaMA-1 and LLaMA-2 show that QLLM is able to obtain accurate quantized models efficiently. For example, QLLM quantizes the 4-bit LLaMA-2-70B within 10 hours on a single A100-80G GPU, outperforming the previous state-of-the-art method by 7.89% on the average accuracy across five zero-shot tasks. Code is available at ZIP Lab and ModelTC.
@@ 5 Experiments
Quantization settings. In alignment with prior research (Dettmers et al., 2022; Shao et al., 2023), we use per-channel weight quantization and per-token activation quantization. Following (Shao et al., 2023; Liu et al., 2023), we quantize all weights and intermediate activations, with the exception of the Softmax output probability, which is maintained at full precision. Following OmniQuant (Shao et al., 2023), we focus on 4- and 6-bit weights and activations quantization. Additionally, we also explore 4-bit weights and 8-bit activations quantization, aiming for hardware-friendly configurations while maintaining high performance. We exclude 8-bit quantization as SmoothQuant (Xiao et al., 2023) is able to achieve lossless performance.
@@
Implementation details. Following OmniQuant (Shao et al., 2023), we construct the calibration set with 128 randomly sampled sequences from WikiText2, each with a sequence length of 2048. QLLM begins by applying channel reassembly prior to all linear projection layers, excluding the attention output projection layer, followed by performing error correction on the resulting model. The rank of the introduced low-rank parameters is set to 4, and these parameters are trained for 10 epochs with a mini-batch size of 1. We carry out the reconstruction using 4 Attention-FFN blocks. AdamW (Loshchilov & Hutter, 2019) with a linear learning rate decay scheduler is used following (Yao et al., 2022). The learning rate is set to in most experiments; for LLaMA-2-70B, it is set to . All training experiments are conducted on a single NVIDIA A100 80G GPU. We use the Language Model Evaluation Harness toolbox (Gao et al., 2021) for evaluation.
@@ 5.1 Main Results
We report the results on LLaMA-1 and LLaMA-2 families in Table 1, and Table A in Appendix. Note that W6A6 has limited hardware support in real-world applications. However, our QLLM still demonstrates performance benefits in these settings, consistently surpassing OmniQuant in terms of lower perplexity across all models on both WikiText2 and C4 and achieving comparable accuracy on 5 zero-shot tasks. Remarkably, with W4A8 quantization, our method incurs only a minimal performance reduction. While the absolute performance gains with 6-bit quantization might seem modest, this is partly due to the less pronounced effect of activation outliers at this bitwidth. When focusing on extremely low-bitwidth quantization (i.e., 4-bit), activation outliers serve as the performance bottleneck, thereby highlighting the importance of suppressing the outliers. In this case, our QLLM achieves significantly higher zero-shot accuracy and much lower perplexity than the contenders. For example, QLLM quantized 4-bit LLaMA-1-65B outperforms OmniQuant counterpart by an average of 3.42% in accuracy across five zero-shot tasks. Remarkably, for LLaMA-7B, our QLLM even surpasses the QAT method, LLM-QAT + SQ, by 8.6% on the average accuracy, which strongly demonstrates the efficacy of our QLLM.
@@ 5.2 Ablation Studies
Effect of efficient gradient-based error correction. After channel reassembly, we implement our QLLM to produce 4-bit LLaMA-7B models with our efficient gradient-based error correction (EEC) and tuning quantized weights directly (TQW) outlined in Section 4.2 to further improve the performance of quantized LLMs and show the results in Table 4. Compared with TQW which tunes all quantized weights, EEC focuses on learning a small set of low-rank weights, which significantly reduces training costs and GPU memory usage while delivering comparable performance. Moreover, the reduced GPU memory demand allows EEC to quantize LLaMA-1-65B on a single 24GB consumer-grade GPU, such as the NVIDIA RTX 4090, a task that is not feasible with TQW. Due to the page limited, we put more results in Section L of the supplementary material.
Inference efficiency. To assess the inference efficiency of our channel reassembly technique, we measure the inference speed of QLLM on NVIDIA RTX 3090 GPUs. We employ W4A4 kernels from QUIK (Ashkboos et al., 2023) codebase. We also conduct a comparative analysis using weight quantization only, utilizing CUDA kernels from AutoGPTQ111https://github.com/PanQiWei/AutoGPTQ. As shown in Table 3, our 4-bit QLLM only incurs 4% additional cost relative to W4A4 but achieves a notable 1.96 speedup over FP16. Notably, our channel reassembly strategy substantially mitigates losses attributed to quantizing outliers (see Table E), with only a slight extra computational overhead. For the detailed inference cost of channel disassembly and assembly, please refer to Section N of the supplementary material.
@@ 6 Conclusion and Future Work
In this paper, we have proposed an accurate and efficient post-training quantization approach for low-bit LLMs, dubbed QLLM. The core of our QLLM lies in a novel adaptive channel reassembly paradigm that effectively addresses activation outliers, a pivotal factor contributing to the performance bottleneck in quantizing LLMs. The key idea involves reallocating outlier magnitudes to other channels, accomplished through a process of channel disassembly followed by assembly. We have further proposed a quantization-aware, parameter-efficient fine-tuning strategy that leverages calibration data to compensate for the information loss resulting from quantization. Extensive experiments on LLaMA model series have demonstrated the promising performance and training efficiency of QLLM. In terms of limitations, our proposed channel reassembly involves introducing additional operations to decompose and aggregate channels during runtime, thereby incurring additional inference costs. A potential solution to improve inference efficiency is to explore kernel fusing (Wang et al., 2010), aiming to fuse disassembly, assembly and layer normalization into a single operator. Another way is to aggregate more similar or unimportant channels (Sun et al., 2023) than those disassembled to achieve higher speedup.
@@ Acknowledgments
We sincerely thank Shenghu Jiang for his help in implementing the efficient Triton kernel.
@@ Appendix E More results on LLaMA-2 family
We provide additional results for the LLaMA-2 family in Table A. The observations from these results are consistent with the phenomena identified in the LLaMA-1 family. Note that the varied performance of OmniQuant on W6A6 and W4A4 for LLaMA-2-70B can be attributed to the architecture of LLaMA-2-70B, which employs grouped-query attention (Ainslie et al., 2023) where each group of queries shares a single key and value head. Such architecture makes the learnable equivalent transformation in OmniQuant incompatible with grouped-query attention. In W6A6 settings, the impact of activation outliers is relatively minor, enabling partial learnable equivalent transformations to suffice in maintaining performance. However, in the W4A4 settings, the effect of activation outliers becomes more prominent. Under these conditions, the partial learnable equivalent transformation is insufficient to address the outlier issue, leading to notably poorer performance. Notably, our QLLM significantly outperforms the state-of-the-art post-training quantization (PTQ) methods, demonstrating a substantial margin of improvement in 4-bit quantization. For example, QLLM quantized 4-bit LLaMA-2-70B outperforms SmoothQuant counterpart by an average of 7.89% on the accuracy, which shows the promising results of our method.
@@ Appendix H More results in terms of channel disassembly only
Appendix I More comparisons with other outlier handling methods
To further show the effectiveness of channel reassembly, we compare our method with previous outlier handling methods which employ gradient-free methods to learn mathematically equivalent transformations. For fair comparisons, we do not apply efficient error correction. From Table E, all methods exhibit comparable performance at 6-bit quantization. However, for 4-bit quantization, channel reassembly significantly surpasses other methods by a large margin, particularly for larger models.
| LLaMA-1-7B | W6A6 | SQ | 76.65 | 53.11 | 40.10 | 71.52 | 61.88 | 60.65 |
| W6A6 | OS+ | 76.82 | 51.35 | 41.13 | 71.42 | 65.98 | 61.34 | |
| W6A6 | CR | 76.88 | 52.31 | 40.87 | 71.37 | 64.33 | 61.15 | |
[…]
| W4A4 | OS+ | 62.73 | 39.98 | 30.29 | 44.39 | 52.96 | 46.07 | |
| W4A4 | CR | 66.92 | 42.55 | 32.34 | 54.31 | 50.04 | 49.23 | |
| LLaMA-1-13B | W6A6 | SQ | 77.80 | 56.36 | 42.58 | 75.11 | 68.11 | 63.99 |
| W6A6 | OS+ | 78.29 | 56.90 | 43.09 | 75.09 | 69.22 | 64.52 | |
| W6A6 | CR | 78.02 | 56.69 | 42.41 | 74.70 | 70.01 | 64.37 | |
[…]
| W4A4 | OS+ | 63.00 | 40.32 | 30.38 | 53.61 | 51.54 | 47.77 | |
| W4A4 | CR | 67.57 | 43.77 | 31.48 | 60.78 | 56.04 | 51.93 |
Appendix J More results in terms of the efficient error correction only
Using EEC only without our channel reassembly results in suboptimal performance as it suffers from activation outlier issues. To demonstrate this, we applied EEC only to quantize LLaMA-1-7B to 4-bit, using the same training settings as our QLLM but with varying numbers of calibration samples. From Table F, even with an increased amount of calibration data, the performance of the EEC only significantly lags behind our QLLM. These results strongly demonstrate the effectiveness of channel reassembly in addressing activation outliers, thereby substantially improving performance.
@@ Appendix I More comparisons with other outlier handling methods
Appendix K More results in terms of tuning quantized weights only
The effectiveness of TQW is highly dependent on our channel reassembly. To demonstrate this, we applied TQW only to quantize LLaMA-1-7B to 4-bit using the same training settings as QLLM and show the results in Table G. The results clearly indicate that the absence of our adaptive channel reassembly results in significantly reduced performance for TQW. This underscores the vital role of channel reassembly in addressing activation outliers and thus improving model performance.
@@
Appendix L More comparisons between efficient error correction and tuning quantized weights directly
To further show the effectiveness of our efficient error correction (EEC), we conduct more comparisons between EEC and tuning quantized weights (TQW) directly on small model and report the results in Table H. The results show that employing EEC not only maintains comparable performance but also markedly improves training speed and significantly reduces GPU memory usage over TQW. It is worth noting that there is a trade-off between GPU memory (i.e., #Attn-FFN blocks) and performance. Leveraging EEC even allows us to perform reconstruction for 16 Attention-FFN blocks simultaneously, thereby significantly improving performance while preserving a similar training speed and a reasonable increase in GPU memory.
@@
Appendix M Effect of the weight merging in efficient error correction
As explained in Section 4.2, for , the low-rank weights and bring not only additional inference overhead due to the matrix multiplication between the full-precision and but also extra storage burden. To address this, we perform weight merging by after the reconstruction, which effectively avoids overhead but introduces additional quantization error. For 4-bit quantization, we empirically observe that merging the low-rank weights into the frozen weights using does not lead to an increase in outliers. This finding is supported by the notably low MSE levels for channel-wise , , and maximum/minimum values before and after the weight merging process across in Table I. Moreover, our weight merging only leads to small quantization error, as shown in Table J. Note that even small deviations can aggregate throughout the network, leading to the performance drop. To address this, as shown in Section 4.2, we further employ sequential reconstruction to mitigate errors from previous layers, resulting in only a negligible performance drop. To demonstrate this, we compare the performance of QLLM with and without the weight merging. From Table K, the weight merging only leads to a slight increase in perplexity.
@@
Appendix N More results regarding inference efficiency
Following (Guo et al., 2020; Wang et al., 2020), we use Bit-Operation (BOP) count to measure the theoretical inference complexity of our QLLM. From Table L, our 8-bit QLLM incurs only a marginal increase in BOPs when compared to the INT8 model but substantially lower than those of the FP16 counterpart, which shows the efficiency of our method.
@@
We further show the inference time of channel disassembly and assembly of our QLLM in Table M. From the results, channel disassembly results in additional inference costs due to the extra channels. These additional channels often don’t align with GPU-friendly multiples like 32 or 64, leading to less efficient GPU use. Using our channel assembly maintains the original channel count, ensuring better GPU utilization and mitigating the extra inference costs from disassembly. As a result, the quantized models with both channel disassembly and assembly achieve higher throughput compared to the ones with disassembly only, which demonstrates the necessity of channel assembly.
@@
Appendix O More results regarding training efficiency
We assess the training efficiency of our method in comparison to OmniQuant on a single NVIDIA A100 80G GPU. The GPU training hours for both methods are presented in Table N. The results reveal that the training cost of our QLLM can be up to 1.93 faster than OmniQuant, showing the exceptional training efficiency of our QLLM.
@@
Appendix P Effect of different calibration sets
We apply QLLM to yield 4-bit LLaMA-7B using different calibration sets and report the results in Table O. From the results, we observe that the choices of calibration set have a minor effect, as the performance remains relatively consistent across different sets. For fair comparisons, following OmniQuant (Shao et al., 2023), we use WikiText2 as a calibration set by default.
@@
Appendix Q Effect of different numbers of calibration samples
We apply QLLM to yield 4-bit LLaMA-7B using different numbers of calibration samples from WikiText2 and show the results in Table P. The results reveal a positive correlation between the performance of the quantized model and the number of calibration samples, indicating that utilizing more samples generally leads to better performance. This trend underscores the importance of the calibration phase, where leveraging a larger sample pool can provide a more comprehensive representation of the data distribution, enabling more accurate quantization. However, it is also imperative to consider the computational and memory overhead associated with an increasing number of calibration samples. There is an inherent trade-off between achieving higher model performance and maintaining computational efficiency.
@@
Appendix R More results about the expansion ratios of the quantized LLM